OpenStudy (anonymous):

can anyone help me get started on this? dh/dt= 0.016(2.5-h) Solve the differential equation to find h in terms of t.

OpenStudy (anonymous):

dh/(2.5-h)=0.016 dt now integrating, -log(2.5-h)=0.016 t+c

OpenStudy (anonymous):

Is this enough, ersp?

OpenStudy (anonymous):

i think so my brains kinda died doing this work XD thanks :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

actually i dont get where the minus log comes from when integrating

OpenStudy (anonymous):

Comes from the -h...

OpenStudy (anonymous):

\[\int\limits_{}{}\frac{dh}{2.5-h}=-\int\limits_{}{}\frac{(2.5-h)'dh}{2.5-h}=-\log (2.5-h)+c\]

OpenStudy (anonymous):

I got bumped. Is this explanation okay?

OpenStudy (anonymous):

i think so cheers

OpenStudy (anonymous):

so what method is this?

OpenStudy (anonymous):

Method? To solve the differential equation?

OpenStudy (anonymous):

The d.e. is separable.

OpenStudy (anonymous):

Or, you can just take the reciprocal of both sides and integrate directly, but it's the same thing.

OpenStudy (anonymous):

thanks makes a bit more sense to me now, my mind is really failing me today!

OpenStudy (anonymous):

np ;)

OpenStudy (anonymous):

shouldnt both sides be logged?

OpenStudy (anonymous):

No. It goes like this:\[\frac{dh}{dt}=0.016(2.5-h) \rightarrow \frac{dh}{2.5-h}=0.016dt \rightarrow \int\limits_{}{}\frac{dh}{2.5-h}=\int\limits_{}{}0.016dt\]\[\rightarrow -\log (2.5-h) = 0.016t + c \rightarrow \log (2.5-h)=-0.016t+c_1 \rightarrow 2.5-h=c_2e^{-0.016t}\]\[\rightarrow h=2.5-ce^{-0.016t}\]

OpenStudy (anonymous):

I added subscripts as I went along, but we just lump constants into constants, so normally, the subscripts are left off...which is what I did in the end.

OpenStudy (anonymous):

ah that makes sense thanks very much, this work ive been throwing my head in to a wall for the last few days!

OpenStudy (anonymous):

lol, it happens :p

OpenStudy (anonymous):

Thanks for fanning me, ersp :D

OpenStudy (anonymous):

after that much help it seems logical! XD

OpenStudy (anonymous):

ah it doesnt seem to fit for the next part of the question, at max h is 2.5, how long does it take to get to this point, so i have to rearrange it to make t the subject but the values dont seem to fit at all!